Solve the following equations:
10p = 10 = 100.......

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

First we have to subtract 10 to the both sides of the equation,
Then, we get,
= 10p + 10 - 10 = 100 - 10
= 10p = 90
Now,
We have to divide both sides of the equation by 10,
Then, we get,
= 10p/10 = 90/10
=p= 9

Thanks
Prachi Singh
2 years ago
Dear Student,
We can solve it as follows :
10p + 10 = 100
10 p = 100-10
10 p = 90
P = 90/10
P = 9
 
Hope it helps!
Thank you